The first term in the sequence is 256. After the first term, each term is multiplied by two and then divided by four. What is the value of the 9th term?
(A) 8 (B) 4 (C) 2 (D) 1 (E) 0

Answer by Edwin McCravy(20056)

The first term in the sequence is 256. After the first term, each term is multiplied by two and then divided by four. What is the value of the 9th term?

It's the same as taking one-half each time.
1. 256
2. 256x2÷4=128
3. 128x2÷4=64
4. 64x2÷4=32
5. 32x2÷4=16
6. 16x2÷4=8
7. 8x2÷4=4
8. 4x2÷4=2
9. 2x2÷4=1
The answer is (D)